# Climbing the Ladder of Interpretability with Counterfactual Concept Bottleneck Models

### v1

@@ 5.3 What if? Counterfactual actionability
CF-CBMs calibrate the variability-plausibility trade-off. (Figure 5)
@@
Another key trade-off in generating counterfactuals corresponds to the tension between producing counterfactuals with high diversity, allowing users to pick among different roadmaps to action, and producing plausible counterfactuals corresponding to user actions that are likely to be feasible in practice. However, learning shortcuts might lead to trivial solutions with a poor variability-plausibility trade-off. For instance, once a plausible counterfactual is found, a trivial solution could be to generate always the same counterfacual, thus leading to minimal diversity, while maximizing the variability increases the chance of generating less plausible counterfactuals. In practice, these two properties are measured by variability and IoU between counterfactuals and ground-truth concept vectors. A higher variability implies that our model can produce a wider range of counterfactuals. A higher IoU value suggests that generated counterfactuals are also representative of the training distribution which contains plausible solutions. Our results show that CF-CBMs balance the optimization of these two properties better than existing methods. The gap w.r.t. BayCon and VCNet is already significant on dSprites where CF-CBMs produce counterfactuals with $+25$ p.p. higher IoU (plausibility) and ×$\times 3$ more variable. In MNIST addition the VAECF’s performances significantly deteriorate, while C-CHVAE mainly struggles in CUB.
CF-CBM (ours) CBM End-to-end NN BayCon C-CHVAE VAE-CF VCNet
What? ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$
Why? ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$
What if? ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$
Concept interventions ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$
Task-driven interventions ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$
Real-time ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$
Table 2: CF-CBMs efficiently address “what?”, “why?”, and “what if?” questions while supporting both concept and task-interventions, as opposed to existing CBMs and counterfactual generators.
CF-CBMs generate accurate task-driven interventions. (Figure 6)
@@
CF-CBMs enable task-driven interventions showing users how to modify concepts to get the desired class label. This represents a novel crucial functionality when users are aware of a desired outcome but lack the knowledge required to intervene on concepts. CF-CBM significantly outperforms other baselines in concept accuracy of task-driven interventions by up to $78$ p.p. Such performance can be explained as the generation of task-driven interventions is conditioned on a user-provided target class label which provides fundamental information for the model to search in the latent concept distribution. Appendix C shows that CF-CBMs attain similar results while applying different levels of noise on predicted concept labels.
CF-CBMs efficiently generate counterfactuals on the fly. (Figure 7)
Generation efficiency is a key feature for real-time applications. In this respect, CF-CBMs significantly outperform iterative models such as C-CHVAE and BayCon. These models quickly struggle in large search spaces involving a high number of concepts: generating a counterfactual takes more than a minute on dSprites (7 concepts) and up to a few hours on CUB (118 concepts). CF-CBMs instead scale well with the size of the data set by requiring less than a second on dSprites and a few seconds on CUB, on par with VAECF and VCNet. This efficiency enables a wider range of applications requiring a high number of concepts and quick feedbacks.
@@ 6 Discussion & Conclusion
CF-CBMs combine features from CBMs and generative models, carrying forward their strengths and limitations. From generative models, CF-CBMs inherit the ability to efficiently approximate the data distribution. However, this approximation does not guarantee counterfactuals’ optimality when compared to exact methods (Wachter et al., 2017). Besides, CF-CBMs require concept annotations to ground explanations, as standard CBMs. Reasoning shortcuts (Marconato et al., 2023), concept impurities (Zarlenga et al., 2023), and information bottlenecks (Espinosa Zarlenga et al., 2022) are also typical limitations of CBM architectures, especially when the concept bottleneck is not complete (Yeh et al., 2020). However, recent works show how CBMs (and the same argument holds for CF-CBMs) could be extended using language models to approximate ground-truth concept annotations (Oikarinen et al., 2023) and concept embedding reasoners to provide robust task explanations reducing information bottlenecks (Barbiero et al., 2023).
Relations with concept-based models.
@@
CF-CBMs share common ground with existing CBMs in modeling concepts in latent spaces. However, current models use the latent concept space to improve task accuracy (Espinosa Zarlenga et al., 2022; Barbiero et al., 2023), promote concept disentanglement (Misino et al., 2022; Marconato et al., 2022), or estimate confidence intervals for CBM’s predictions (Kim et al., 2023). Differently, CF-CBMs leverage the latent concept space to extend the capabilities of the CBM family by introducing two key functionalities (Table 2): counterfactual explanations and task-driven interventions.
Relations with generative structured models.
@@
CF-CBMs are latent generative models characterized by a structured latent space. This characterization aligns them with other methodologies utilizing hierarchical priors (Sønderby et al. (2016); Maaløe et al. (2019); Vahdat and Kautz (2020); Klushyn et al. (2019)), mixture priors (Bauer and Mnih (2019)), and autoregressive priors (Chen et al. (2017); van den Oord et al. (2017); Razavi et al. (2019)). However, the structure of CF-CBMs’ latent variables explicitly captures both the concept and the counterfactual latent space, thus promoting structural interpretability, unlike existing structured latent variable models.
Relation with counterfactual models.
@@
Existing DL methods generating counterfactuals typically offer post-hoc explanations (Goyal et al., 2019; Ghandeharioun et al., 2021; Abid et al., 2022) or require interpretable input features (Guyomard et al., 2023; Pawelczyk et al., 2020; Mahajan et al., 2020). However, applying post-hoc methods directly to unstructured data types (e.g., pixels rather than concepts) significantly reduces counterfactuals’ interpretability (Kim et al., 2018). In contrast, CF-CBMs: (i) are data-agnostic, allowing for the extraction of interpretable counterfactuals regardless of the data type, (ii) provide answers to “why?” questions by design via concept-based explanations, and (iii) support both concept and task-driven interventions (Table 2). Moreover, generating counterfactuals at the concept level brings significant advantages. First, concept-based counterfactuals are more interpretable and actionable, involving changes in high-level semantic units of information rather than raw features. As a side effect, concept-based counterfactuals may also introduce benefits in terms of privacy as sensitive and unique details embedded in input data (e.g., the image of a radiography) remain concealed, thus mitigating the risk of data extraction through repeated model queries (Goethals et al., 2023; Naretto et al., 2022; Pawelczyk et al., 2022).
Broader Impact
@@
The ability to cover the ladder of interpretability represents a fundamental step to calibrate human trust and enhance human-machine interactions. In this work, we show how patching existing interpretable models (such as CBMs) and post-hoc counterfactual generators may represent a possible solution to address “what?”, “why?”, and “what if?” queries, thus covering the three rungs of the “ladder”. However, experimental results show that ad-hoc methods—designed and trained to address these three fundamental queries, such as CF-CBMs—may generate explanations and counterfactuals with higher quality. Besides, the ability to perform task-driven interventions may provide actionable insights to improve CF-CBMs’ predictions whenever users observe a mispredicted class label but do not know how to intervene on concepts. Finally, the proposed latent variable approach enables to efficiently model complex concept dependencies while generating multiple counterfactuals for a given input. This represents a key advantage w.r.t. iterative methods in practical applications where proposing a variety of interpretable counterfactuals on the fly represents a fundamental advantage, as in suggesting the quickest ways home to an autonomous vehicle or developing different treatment strategies under varying clinical scenarios.
7 Acknowledgments
This study was funded by TRUST-ME (project 205121L_214991), SmartCHANGE (GA No. 101080965) and XAI-PAC (PZ00P2_216405) projects.
@@ Appendix A Ammortized inference and ELBO Derivation
Appendix B Experimental details
B.1 Data & task setup
@@
In our experiments we use three different datasets commonly used to evaluate CBMs:
@@
  • dSprites [Matthey et al., 2017] — It is composed of images of one of three objects (square, ellipse, heart) in different positions and of different sizes. Starting from it, we design a binary task, which is to predict if in the image there is at least a square or a heart. We also combine initial images to obtain new ones with more than one object. The concepts for this task are: (1) the presence of a square, (2) the presence of an ellipse, (3) the presence of a heart, (4) the presence of two objects, (5) objects are red, (6) objects are green, (7) objects are blu.
@@
  • MNIST addition — The task of this dataset is to predict the sum of two MNIST digits [Deng, 2012]. The concepts of this task are the one-hot encoding of the first and the second digits concatenated together.
@@
  • CUB [Welinder et al., 2010] — It contains pictures of birds and the final task is to predict their species (among 200 species). The concepts are 118 bird features that human annotators selected.
B.2 Evaluation Metrics

### v3

@@ 5.3 Why not? Counterfactual actionability
Figure 3: CF-CBMs balance the trade off between counterfactuals’ reliability (proximity) and user effort ($\Delta$-Sparsity). The arrow ()($\rightarrow$)( → ) points towards optimal values. Pareto-optimal models form the frontier of the shaded region, whereas dominated solutions are located within the shaded region.
Figure 4: CF-CBMs calibrate the trade off between counterfactuals’ diversity (variability) and plausibility (IoU). The arrow ()($\rightarrow$)( → ) points towards optimal values. Pareto-optimal models form the frontier of the shaded region, whereas dominated solutions are located within the shaded region. “CF-CBM $n$” samples $n$ counterfactuals from the latent distribution as described in App. B.
Figure 5: CF-CBMs generate reliable (proximity) and accurate interventions (ROC AUC Int.). The arrow ()($\rightarrow$)( → ) points towards optimal values. Pareto-optimal models form the frontier of the shaded region, whereas dominated solutions are located within the shaded region.
@@
CF-CBMs calibrate the variability-plausibility trade-off. (Figure 4) Another key trade-off in generating counterfactuals corresponds to the tension between producing counterfactuals with high diversity, allowing users to pick among different roadmaps to action, and producing plausible counterfactuals corresponding to user actions that are likely to be feasible in practice. However, learning shortcuts might lead to trivial solutions with a poor variability-plausibility trade-off. For instance, once a plausible counterfactual is found, a trivial solution could be to generate always the same counterfacual, thus leading to minimal diversity, while maximizing the variability increases the chance of generating less plausible counterfactuals. In practice, these two properties are measured by variability and IoU between counterfactuals and ground-truth concept vectors. A higher variability implies that our model can produce a wider range of counterfactuals. A higher IoU value suggests that generated counterfactuals are also representative of the training distribution which contains plausible solutions. Our results show that CF-CBMs balance the optimization of these two properties better than existing methods. The gap w.r.t. BayCon and VCNet is already significant on dSprites where CF-CBMs produce counterfactuals with $+25$ p.p. higher IoU (plausibility) and ×$\times 3$ more variable. In MNIST addition the VAECF’s performances significantly deteriorate, while C-CHVAE mainly struggles in CUB. On CIFAR10, all baselines perform poorly on the IoU metric, while on the SIIM Pneumothorax dataset, they also struggle with variability, with the exception of C-CHVAE.
@@
CF-CBMs generate accurate task-driven interventions. (Figure 5) CF-CBMs enable task-driven interventions, showing users how to modify concepts to get the desired class label. This represents a novel crucial feature when users know the desired outcome but lack the knowledge to intervene on concepts. CF-CBM significantly outperforms other baselines in concept accuracy of task-driven interventions by up to $78$ p.p. Such performance can be explained as the generation of task-driven interventions is conditioned on a user-provided target class label which provides fundamental information for the model to search in the latent concept distribution. Appendix D shows that CF-CBMs attain similar results while applying different levels of noise on predicted concept labels.
@@ 6 Discussion & Conclusion
Limitations. CF-CBMs combine features from CBMs and generative models, carrying forward their strengths and limitations. From generative models, CF-CBMs inherit the ability to efficiently approximate the data distribution. However, this approximation does not guarantee counterfactuals’ optimality when compared to exact methods (Wachter et al., 2017). Besides, CF-CBMs require concept annotators (humans or machines (Oikarinen et al., 2023)) to ground explanations, as standard CBMs. Reasoning shortcuts (Marconato et al., 2023), concept impurities (Zarlenga et al., 2023), and information bottlenecks (Espinosa Zarlenga et al., 2022) are also typical limitations of CBM architectures, especially when the concept bottleneck is not complete (Yeh et al., 2020).
Table 4: CF-CBMs efficiently address “what?”, “how?”, and “why not?” questions while supporting both concept and task-interventions, as opposed to existing CBMs and counterfactual generators.
CF-CBM (ours) CBM Black Box BayCon C-CHVAE VAE-CF VCNet
What? ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$
How? ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$
Why not? ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$
Concept interventions ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$
Task-driven interventions ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$
Real-time ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\bm{\checkmark}}$
@@
Relations with concept-based models. CF-CBMs share common ground with existing CBMs in modeling concepts in latent spaces. However, current models use the latent concept space to improve task accuracy (Espinosa Zarlenga et al., 2022; Barbiero et al., 2023; Debot et al., 2024), promote concept disentanglement (Misino et al., 2022; Marconato et al., 2022), model concept interactions (Xu et al., 2024; Barbiero et al., 2024), or estimate task confidence intervals (Kim et al., 2023). Differently, CF-CBMs leverage the latent concept space to extend CBMs’ capabilities by introducing two key functionalities (Table 4): counterfactual explanations and task-driven interventions.
@@
Relations with generative structured models. CF-CBMs are latent generative models characterized by a structured latent space. This characterization aligns them with other methodologies utilizing hierarchical priors (Sønderby et al. (2016); Maaløe et al. (2019); Vahdat & Kautz (2020); Klushyn et al. (2019)), mixture priors (Bauer & Mnih (2019)), and autoregressive priors (Chen et al. (2017); van den Oord et al. (2017); Razavi et al. (2019)). However, the structure of CF-CBMs’ latent variables explicitly captures both the concept and the counterfactual latent space, thus promoting structural interpretability, unlike existing structured latent variable models.
@@
Relation with counterfactual models. Existing DL methods generating counterfactuals typically offer post-hoc explanations (Ghandeharioun et al., 2021; Abid et al., 2022) or require interpretable input features (Guyomard et al., 2023; Pawelczyk et al., 2020; Mahajan et al., 2020). However, applying post-hoc methods to unstructured data (e.g., pixels instead of concepts) significantly reduces counterfactuals’ interpretability (Kim et al., 2018). In contrast, CF-CBMs: (i) are data-agnostic, extracting interpretable counterfactuals regardless of data type, (ii) provide answers to “how?” questions by allowing interventions, (iii) modify the decision-making process by optimizing counterfactuals through joint optimization leading to relying on fewer concepts and (iv) support both concept and task-driven interventions (Table 4). As a side effect, concept-based counterfactuals may also introduce benefits in terms of privacy as sensitive and unique details embedded in input data (e.g., the image of a radiography) remain concealed, thus mitigating the risk of data extraction through repeated model queries (Goethals et al., 2023; Naretto et al., 2022; Pawelczyk et al., 2022).
@@
Broader Impact. The ability to answer all the three questions represents a fundamental step to calibrate human trust and enhance human-machine interactions. In this work, we show how combining existing interpretable models (such as CBMs) and post-hoc counterfactual generators may represent a possible solution to address “what?”, “how?”, and “why not?” queries. Experimental results show that training the counterfactual generator jointly with the CBM leads to two key improvements: (i) it alters the model’s decision-making process, making the model rely on fewer important concepts (leading to simpler explanations), and (ii) it significantly increases the causal effect of concept interventions on class predictions, making the model more responsive to these changes. Besides, the ability to perform task-driven interventions may provide actionable insights to improve CF-CBMs’ predictions whenever users observe a mispredicted class label but do not know how to intervene on concepts. Finally, the proposed latent variable approach enables to efficiently model concept dependencies while generating multiple counterfactuals for a given input. This represents a key advantage w.r.t. iterative methods in practical applications where proposing a variety of interpretable counterfactuals on the fly represents a fundamental advantage.
Acknowledgments
GD acknowledges support from the European Union’s Horizon Europe project SmartCHANGE (No. 101080965). PB acknowledges support from Swiss National Science Foundation projects TRUST-ME (No. 205121L_214991) and IMAGINE (No. TMPFP2_224226). FG has been supported by the Partnership Extended PE00000013 - “FAIR - Future Artificial Intelligence Research” - Spoke 1 “Human-centered AI”. MG acknowledges support from Swiss National Science Foundation projects XAI-PAC (No. PZ00P2_216405). GM acknowledges support from the KU Leuven Research Fund (STG/22/021, CELSA/24/008) and from the Flemish Government under the "Onderzoeksprogramma Artificiële Intelligentie (AI) Vlaanderen" programme.
@@ Appendix A Ammortized inference and ELBO Derivation
Appendix B Take the best bet or contemplate a multiverse?
Appendix C Experimental details
C.1 Data & task setup
@@
In our experiments we use five different datasets commonly used to evaluate CBMs:
@@
  • dSprites (Matthey et al., 2017) — It is composed of images of one of three objects (square, ellipse, heart) in different positions and of different sizes. Starting from it, we design a binary task, which is to predict if in the image there is at least a square or a heart. We also combine initial images to obtain new ones with more than one object. The concepts for this task are: (1) the presence of a square, (2) the presence of an ellipse, (3) the presence of a heart, (4) the presence of two objects, (5) objects are red, (6) objects are green, (7) objects are blue.
@@
  • MNIST addition — The task of this dataset is to predict the sum of two MNIST digits (Deng, 2012). The concepts of this task are the one-hot encoding of the first and the second digits concatenated together.
@@
  • CUB (Welinder et al., 2010) — It contains pictures of birds and the final task is to predict their species (among 200 species). The concepts are 118 bird features that human annotators selected.
  • CIFAR10(Krizhevsky et al., ) — This dataset contains images of objects, with the task of classifying them into one of 10 classes. Concepts are automatically extracted using the method proposed in Oikarinen et al. (2023), resulting in 143 concepts.
  • SIIM Pneumothorax(Zawacki et al., 2019) — This dataset contains X-ray scans of lungs, with the task of determining whether the patient has pneumothorax. Concepts are automatically extracted following Oikarinen et al. (2023), leveraging CXR-CLIP, resulting in 19 concepts.
C.2 Evaluation Metrics